Stearin, also known as tristearin, is a simple triglyceride formed by three chains of stearic acid. Its biochemical conversion (reaction of stearin in the presence of oxygen) into carbon dioxide and water produces energy.

a) Write a balanced equation for the conversion of stearin into carbon dioxide and water. (20%)

b) Calculate the bond energy for the C-O chemical bond knowing that the enthalpy of the reaction is -26640.5 kJ mol-1 and knowing the following bond energies: C-C (348 kJ/mol), C-H (412 kJ/mol), C=O (740 kJ/mol), O=O (497 kJ/mol) and O-H (463 kJ/mol). (80%)
1
Expert's answer
2019-04-15T07:55:28-0400

Stearin, also known as tristearin, is a simple triglyceride formed by three chains of stearic acid. Its biochemical conversion (reaction of stearin in the presence of oxygen) into carbon dioxide and water produces energy.


a) Write a balanced equation for the conversion of stearin into carbon dioxide and water. (20%)

b) Calculate the bond energy for the C-O chemical bond knowing that the enthalpy of the reaction is -26640.5 kJ mol-1 and knowing the following bond energies: C-C (348 kJ/mol), C-H (412 kJ/mol), C=O (740 kJ/mol), O=O (497 kJ/mol) and O-H (463 kJ/mol). (80%) 

Solution:

a)    2C57H110O6 + 163 O2   ->   114 CO2 + 110 H2O

b)     ΔHrxn = -26640.5 kJ/mol


C-O - ? kJ/mol